# FireFly-S: Exploiting Dual-Side Sparsity for Spiking Neural Networks Acceleration with Reconfigurable Spatial Architecture

### v2

@@ Abstract
Spiking Neural Networks (SNNs), with their brain-inspired structure using discrete spikes instead of continuous activations, are gaining attention for their potential of efficient processing on neuromorphic chips. While current SNN hardware accelerators often prioritize temporal spike sparsity, exploiting sparse synaptic weights offers significant untapped potential for even greater efficiency. To address this, we propose FireFly-S, a Sparse extension of the FireFly series. This co-optimized software-hardware design focusing on leveraging dual-side sparsity for acceleration. On the software side, we propose a novel algorithmic optimization framework that combines gradient rewiring for pruning and modified Learned Step Size Quantization (LSQ) tailored for SNNs, which achieves remarkable weight sparsity exceeding 85% and enables efficient 4-bit quantization with negligible accuracy loss. On the hardware side, we present an efficient dual-side sparsity detector employing a Bitmap-based sparse decoding logic to pinpoint the positions of non-zero weights and input spikes. The logic allows for the direct bypassing of redundant computations, thereby enhancing computational efficiency. Different from the overlay architecture adopted by previous FireFly series, we adopt a parametric spatial architecture with inter-layer pipelining that can fully exploit the fine-grained programmability and reconfigurability of Field-Programmable Gate Arrays (FPGAs), enabling fast deployment for various models. A spatial-temporal dataflow is also proposed to support such inter-layer pipelining and avoid long-term temporal dependencies. In experiments conducted on the MNIST, DVS-Gesture and CIFAR-10 datasets, the FireFly-S model achieves 85-95% sparsity with 4-bit quantization and the hardware accelerator effectively leverages the dual-side sparsity, delivering outstanding performance metrics of 10,047 FPS/W on MNIST, 3,683 FPS/W on DVS-Gesture, and 2,327 FPS/W on CIFAR-10.
Index Terms:
Spiking neural networks, field-programmable gate array, hardware accelerator, dual-side sparsity, software-hardware co-design, inter-layer pipelining
@@ I Introduction
Despite the parallel processing capabilities of graphics processing units (GPUs), which are frequently used to deploy SNNs, their architecture is not tailored to the unique characteristics of SNNs. Specifically, GPUs lack consideration for event-driven processing and are not optimized for the 1-bit arithmetic operations commonly found in SNNs, highlighting the need for dedicated hardware solutions. In response to these shortcomings, researchers have explored the use of application-specific integrated circuits (ASICs) and field-programmable gate arrays (FPGAs). Notable developments include IBM’s TrueNorth chip [3] and Intel’s Loihi [4] and its successor Loihi2 [5], which are ASICs designed explicitly for SNNs and demonstrate marked improvements in energy efficiency and event-based processing. FPGAs also play a crucial role in SNNs acceleration, offering a flexible platform for exploring various optimization strategies. Some designs leverage this flexibility to enhance both arithmetic and memory efficency, as in our previous work, FireFly [6] and FireFly v2 [7], while others focus on exploiting the sparse nature of spikes through event-driven processing [8] [9].
However, although these designs have made significant strides in improving the performance and energy efficiency of SNN hardware accelerators, they often overlook the potential for further optimization in areas such as network topology and synaptic connectivity. Exploiting the dual-side sparsity, specifically synaptic weight sparsity and activation spike sparsity, can substantially reduce both computational load and network parameters. Moreover, their reliance on high-precision data types like FP32, FP16 for weights and membrane potentials [9] [10], also increases the demand for DSP resources and large memory footprints, hindering deployment and limiting efficency on resource-constrained platforms.
@@
    FireFly-S leverages a fully unfolded spatial architecture with inter-layer pipelining, eliminating off-chip memory access and concurrently processing multiple layers. We also presents a inter-layer dataflow tailored for SNNs, along with a decicated logic to bridge the dataflow between layers, minimizing the need for extensive on-chip buffering.
The remaining sections of the paper are organized as follows: Section II reviews the background on SNNs compression and acceleration. Section III details the algorithmic optimization framework proposed in FireFly-S. Section IV describes the hardware acceleration strategy, encompassing both dataflow and sparse representation. Section V presents the hardware micro architecture of FireFly-S, featuring a sparse processor and dataflow orchestrator. Section VI provides experimental results and evaluations. Section VII concludes the paper and discusses future work.
@@ II-A Spiking Neuron Model and Dynamic Equation
Equation (2) accumulates the input current $I_{i}(t)$ from all presynaptic neurons $j$ with weights $w_{ij}$ and spikes $s_{j}(t)$, as well as a bias term $b_{i}$. Equation (3) updates the membrane potential $V_{i}(t)$ based on the input current and the membrane time constant. As for IF neurons, this decay process is absent, and the membrane potential simply integrates the input current without any leakage. Equation (4) checks if the membrane potential exceeds the threshold $V_{th}$, emits a spike if true and resets the potential to $V_{reset}$. In pratical implementations, the membrane time constant $\tau_{m}$ is often set to 2, and the potential reseting value $V_{reset}$ is set to 0, with the membrane resistance $R_{m}$ normalized to 1.
@@ II-B Spiking Neural Networks Compression
However, while pruning can induce dual-side sparsity in both weights and activations, and quantization can further boost the compression ratio, there is limited work on effectively combining these two techniques during training, especially in a hardware-friendly manner that considers the efficency constraints presented by neuromorphic hardware for SNNs, such as fully integer arithmetic within neuron computations.
@@ II-C Spiking Neural Networks Accelerators
Hardware accelerators for SNNs can be broadly categorized into two approaches: general-purpose optimization and specialized acceleration. General-purpose methods focus on optimizing hardware architecture or modules for broader applicability. In contrast, specialized accelerators exploit the unique characteristics of SNNs, such as sparsity and event-driven dynamics, to achieve higher efficiency. For instance, Li et al. presented FireFly [6] and FireFly-v2 [7], which leverage the DSP48E2 units in Xilinx FPGAs and a spatial-temporal dataflow to optimize the accelerator for overlay architectures. Targeting event-driven characteristics, Panchapakesan et al. proposed SyncNN [8], which encodes only activated neurons and quantizes LeNet to 4 bits. Chen et al. introduced SiBrain [10], featuring a sparse detection and response unit based on a channel-cache and blockmultiplex design.
@@ III-A Compression Sequence
The sequence of pruning and quantization shows impact on the performance of neural networks. Some studies advocate for a quantize-then-prune approach [22] [23], while others, such as Zhang et al. [24], suggest that the optimal sequence may depend on the network architecture. This indicates that different networks exhibit varying tendencies regarding the order of compression techniques. Therefore, treating pruning and quantization as separate steps may lead to suboptimal performance across various network, and any necessary post-processing or recalibration [25] [26] to maintain accuracy can become complex and challenging.
@@ III-B Compression Methodologies
The pruning algorithm we adopt is based on gradient rewiring [19]. In this method, each synaptic weight $w$ is represented as $w=s\odot\text{ReLU}(\theta)$, where $\theta$ represents the synaptic parameter, and $\odot$ denotes the element-wise product. The sign factor $s$, determined by the sign of the initial weight, indicates whether the connection is excitatory or inhibitory. This representation is illustrated in the left part of Fig. 2. During training, $\theta$ is influenced by a Laplace prior distribution to impose a regulartization effect. Specifically, a penalty term $-\alpha\,\text{sign}(\theta-\mu)$ is applied, where $\alpha$ and $\mu$ represent the scale and location parameters, respectively[19].This penalty pushes some $\theta$ values towards non-positive numbers, thereby eliminating the corresponding connections.
Our approach builds on this by applying weight decay to $\theta$, scaling $\theta$ by $(1-\eta\lambda)$ during each update, where $\eta$ is the learning rate and $\lambda$ as the weight decay factor. This introduces an $\ell_{2}$ regularization effect, which not only encourages more $\theta$ values to approach zero, promoting sparsity, but also helps avoid sharper minima and improves generalization during optimization [29] [30]. Besides, to enable more comprehensive network compression, we apply this pruning strategy to both convolutional and fully connected layers, including bias parameters that optimized by AdamW [31].
@@
Output: List of active channels $A$
@@ 9                   $A\leftarrow A\cup\{c_{o}\}$;
[…]
@@
These two key distinctions also offer two major benefits: 1) Learnable per-channel scales enable the model to capture and address quantization errors arising from various computations during training, including neuron dynamics like the decay process. This results in improved accuracy, as demonstrated in Section VI. This end-to-end training process ensures that, once trained, the model can be deployed directly without any additional precision loss. In contrast, statistical methods require a post-training process, and since the scales are based on weight statistics, applying them directly to biases and neuronal thresholds can cause mismatches and lead to accuracy degradation. 2) By sharing the quantization scale across weights within the same channel, along with the corresponding bias and neuronal threshold, our approach ensures consistent scaling across parameters. This eliminates the need for rescaling during weight accumulation, bias addition, and threshold comparison, enabling efficient integer-only inference. Additionally, due to the unique properties of spiking neurons, there is no need to rescale per-channel activations during inference compared to ANNs quantization, as the per-channel quantization scales are absorbed into the membrane thresholds. Based on these characteristics, our hardware accelerator can perform all computations using integer arithmetic, simplifying the computational logic and memory footprint.
@@ IV-A Dataflow and Parallelism Schemes
However, a key distinction arises between FireFly-S and its predecessors. While both FireFly and FireFly v2 utilize overlay architectures, relying on off-chip memory for feature map and weight storage, FireFly-S adopts a spatial architecture to enable full inter-layer pipelining. This shift necessitates a departure from FireFly v2’s dataflow strategy. In detail, solutions are required to address the two main challenges posed by spatial architectures: 1) the tight on-chip memory constraints, which demand efficient memory utilization, and 2) the need for a consistent dataflow pattern to ensure efficient processing across layers and avoid pipeline stalls.
Directly applying FireFly v2’s dataflow to a spatial architecture would prove inefficient. FireFly v2, designed with off-chip memory reliance, employs an output-channel outermost loop. This approach, while functional, leads to excessive on-chip memory consumption in a spatial architecture. Computing each output channel requires access to all input feature maps, necessitating storage of all layer feature maps on-chip, amounting to $L\times F_{h_{o}}\times F_{w_{o}}\times T\times C_{i}$ memory, where $L$ represents the number of layers. Such demands would deteriorate the limits of on-chip memory.
Besides, the output channel in this layer is the input channel in the next layer, so it is chanllenging to design a consistent dataflow pattern. But we can minimize the storage and data movement. Employing an output-channel outermost loop and an input-channel innermost loop, as in FireFly v2, would lead to significant pipeline stalls in a spatial architecture. Each layer would need to wait for all output channel computations of the previous layer to complete, so it could not match the need of non-stall pipeline for spatial architecture.
@@ IV-B Sparse Representation Analysis
This analysis, conducted within a specific experimental environment, simulating a convolutional operation with parameters $(T,C_{o},C_{i},F_{h_{o}},F_{w_{o}},K_{h},K_{w})=(4,64,32,16,16,3,3)$. The setup represents a single layer within the 9-layer network for CIFAR-10 recognition discussed in Section VI, and the selection doesn’t affect the overall trend observed. Fig. 4 illustrates the throughput-to-storage ratio across various sparsity levels, comparing the performance of COO, CSR, and Bitmap sparse representations against a dense representation. While COO and CSR outperform Bitmap at extremely high sparsity levels (above 95%) due to Bitmap’s larger memory footprint, Bitmap consistently demonstrates superior performance at other sparsity levels. This advantage is further amplified with larger vector pair lengths, extending its applicability to a wider range of sparsity. Notably, in Figure 4, the sparsity level is shared by both synaptic weights and activation spikes. Experiments indicate that the sparsity of activation spikes ranges from 70% to 90%. Therefore, the sparsity of the weights must be significantly higher to reach the threshold where Bitmap becomes less efficient than COO or CSR. Given these observations, we adopt the Bitmap representation in FireFly-S, prioritizing its greater generality and higher throughput-to-storage ratio across a wide range of sparsity levels.
@@ V Hardware Micro Architecture
The hardware microarchitecture of the FireFly-S is intricately designed to harness the dual-side sparsity characteristic of SNNs. It features a spatial layout that incorporates inter-layer pipelining, as illustrated in Fig. 5. Each pipeline stage is specifically designed to handle a single layer of inference within the FPGA chip, equipped with on-chip parameters and specialized modules including Padding, Dataflow Orchestrator, Dual-side Sparsity Detector, and Maxpooling.
@@
The Padding module dynamically adjusts input dimensions for convolutional layers, adding padding in ”same” mode to match output and input sizes, while bypassing padding for ”valid” mode or fully connected layers. The Dataflow Orchestrator is pivotal in ensuring consistent dataflow between layers to maintain pipeline efficiency. It also efficiently manages the im2col operation for SNNs, essential for reorganizing input data into a format suitable for convolution operations, all while maintaining a membrane potential storage-free approach. The Dual-side Sparsity Detector module optimizes computations by identifying matching nonzero elements in both activation spikes and synaptic weights, enabling computational savings by bypassing unnecessary operations. Then, it generates specific read commands for relevant data (weights, biases, etc.) and efficiently updates neuron dynamics in the IF/LIF neuron model. Finally, the optional Maxpooling module provides spatial downsampling of the output, the output stream is then pipelined and fed into the next layer for further processing.

### v3

@@ Abstract
Spiking Neural Networks (SNNs), with their brain-inspired structure using discrete spikes instead of continuous activations, are gaining attention for their potential of efficient processing on neuromorphic chips. While current SNN hardware accelerators often prioritize temporal spike sparsity, exploiting sparse synaptic weights offers significant untapped potential for even greater efficiency. To address this, we propose FireFly-S, a Sparse extension of the FireFly series. This co-optimized software-hardware design focuses on leveraging dual-side sparsity for acceleration. On the software side, we propose a novel algorithmic optimization framework that combines gradient rewiring for pruning and modified Learned Step Size Quantization (LSQ) tailored for SNNs, achieving a remarkable weight sparsity exceeding 85% and enabling efficient 4-bit quantization with negligible accuracy loss. On the hardware side, we present an efficient dual-side sparsity detector employing a Bitmap-based sparse decoding logic to pinpoint the positions of non-zero weights and input spikes. The logic allows for direct bypassing of redundant computations, thereby enhancing computational efficiency. Different from the overlay architecture adopted by previous FireFly series, we adopt a parametric spatial architecture with inter-layer pipelining that can fully exploit the fine-grained programmability and reconfigurability of Field-Programmable Gate Arrays (FPGAs), enabling fast deployment for various models. A spatial-temporal dataflow is also proposed to support such inter-layer pipelining and avoid long-term temporal dependencies. In experiments conducted on the MNIST, DVS-Gesture and CIFAR-10 datasets, the FireFly-S model achieves 85–95% sparsity with 4-bit quantization and the hardware accelerator effectively leverages the dual-side sparsity, delivering outstanding performance metrics of 10,047 FPS/W on MNIST, 3,683 FPS/W on DVS-Gesture, and 2,327 FPS/W on CIFAR-10.
@@ I Introduction
Despite the parallel processing capabilities of graphics processing units (GPUs), which are frequently used to deploy SNNs, their architecture is not tailored to the unique characteristics of SNNs. Specifically, GPUs are not designed for event-driven processing and are not optimized for the 1-bit arithmetic operations commonly found in SNNs, highlighting the need for dedicated hardware solutions. In response to these shortcomings, researchers have explored the use of application-specific integrated circuits (ASICs) and field-programmable gate arrays (FPGAs). Notable developments include IBM’s TrueNorth chip [3] and Intel’s Loihi [4] and its successor Loihi2 [5], which are ASICs designed explicitly for SNNs and demonstrate marked improvements in energy efficiency and event-based processing. FPGAs also play a crucial role in SNNs acceleration, offering a flexible platform for exploring various optimization strategies. Some designs leverage this flexibility to enhance both arithmetic and memory efficiency, as in our previous work, FireFly [6] and FireFly v2 [7], while others focus on exploiting the sparse nature of spikes through event-driven processing [8] [9].
However, although these designs have made significant strides in improving the performance and energy efficiency of SNN hardware accelerators, they often overlook the potential for further optimization in areas such as network topology and synaptic connectivity. Exploiting the dual-side sparsity, specifically synaptic weight sparsity and activation spike sparsity, can substantially reduce both computational load and network parameters. Moreover, their reliance on high-precision data types like FP32, FP16 for weights and membrane potentials [9] [10], also increases the demand for DSP resources and large memory footprints, hindering deployment and limiting efficiency on resource-constrained platforms.
@@
    FireFly-S leverages a fully unfolded spatial architecture with inter-layer pipelining, eliminating off-chip memory access and concurrently processing multiple layers. We also present an inter-layer dataflow tailored for SNNs, along with a dedicated logic to bridge the dataflow between layers, minimizing the need for extensive on-chip buffering.
The remaining sections of the paper are organized as follows: Section II reviews the background on SNN compression and acceleration. Section III details the algorithmic optimization framework proposed in FireFly-S. Section IV describes the hardware acceleration strategy, encompassing both dataflow and sparse representation. Section V presents the hardware micro architecture of FireFly-S, featuring a sparse processor and dataflow orchestrator. Section VI provides experimental results and evaluations. Section VII concludes the paper and discusses future work.
@@ II-A Spiking Neuron Model and Dynamic Equation
Equation (2) accumulates the input current $I_{i}(t)$ from all presynaptic neurons $j$ with weights $w_{ij}$ and spikes $s_{j}(t)$, as well as a bias term $b_{i}$. Equation (3) updates the membrane potential $V_{i}(t)$ based on the input current and the membrane time constant. As for IF neurons, this decay process is absent, and the membrane potential simply integrates the input current without any leakage. Equation (4) checks if the membrane potential exceeds the threshold $V_{th}$, emits a spike if true and resets the potential to $V_{reset}$. In practical implementations, the membrane time constant $\tau_{m}$ is often set to 2, and the potential resetting value $V_{reset}$ is set to 0, with the membrane resistance $R_{m}$ normalized to 1.
@@ II-B Spiking Neural Networks Compression
However, while pruning can induce dual-side sparsity in both weights and activations, and quantization can further boost the compression ratio, there is limited work on effectively combining these two techniques during training, especially in a hardware-friendly manner that considers the efficiency constraints presented by neuromorphic hardware for SNNs, such as fully integer arithmetic within neuron computations.
@@ II-C Spiking Neural Networks Accelerators
Hardware accelerators for SNNs can be broadly categorized into two approaches: general-purpose optimization and specialized acceleration. General-purpose methods focus on optimizing hardware architecture or modules for broader applicability. In contrast, specialized accelerators exploit the unique characteristics of SNNs, such as sparsity and event-driven dynamics, to achieve higher efficiency. For instance, Li et al. presented FireFly [6] and FireFly v2 [7], which leverage the DSP48E2 units in Xilinx FPGAs and a spatial-temporal dataflow to optimize the accelerator for overlay architectures. Targeting event-driven characteristics, Panchapakesan et al. proposed SyncNN [8], which encodes only activated neurons and quantizes LeNet to 4 bits. Chen et al. introduced SiBrain [10], featuring a sparse detection and response unit based on a channel-cache and block-multiplex design.
@@ III-A Compression Sequence
The sequence of pruning and quantization shows impact on the performance of neural networks. Some studies advocate for a quantize-then-prune approach [22] [23], while others, such as Zhang et al. [24], suggest that the optimal sequence may depend on the network architecture. This indicates that different networks exhibit varying tendencies regarding the order of compression techniques. Therefore, treating pruning and quantization as separate steps may lead to suboptimal performance across various networks, and any necessary post-processing or recalibration [25] [26] to maintain accuracy can become complex and challenging.
@@ III-B Compression Methodologies
The pruning algorithm we adopt is based on gradient rewiring [19]. In this method, each synaptic weight $w$ is represented as $w=s\odot\text{ReLU}(\theta)$, where $\theta$ represents the synaptic parameter, and $\odot$ denotes element-wise product. The sign factor $s$, determined by the sign of the initial weight, indicates whether the connection is excitatory or inhibitory. This representation is illustrated in the left part of Fig. 2. During training, $\theta$ is influenced by a Laplace prior distribution to impose a regularization effect. Specifically, a penalty term $\alpha\,\text{sign}(\theta-\mu)$ is applied, where $\alpha$ and $\mu$ represent the scale and location parameters, respectively[19]. This penalty pushes some $\theta$ values towards non-positive numbers, thereby eliminating the corresponding connections.
Our approach builds on this by applying weight decay to $\theta$, scaling $\theta$ by $(1-\eta\lambda)$ during each update, where $\eta$ is the learning rate and $\lambda$ as the weight decay factor. This introduces an $\ell_{2}$ regularization effect, which not only encourages more $\theta$ values to approach zero, promoting sparsity, but also helps avoid sharper minima and improves generalization during optimization [29] [30]. Besides, to enable more comprehensive network compression, we apply this pruning strategy to both convolutional and fully connected layers, including bias parameters that are optimized by AdamW [31].
@@
Output: List of active channels AA
@@ 9       $A\leftarrow A\cup\{c_{o}\}$;
Here, $r$ represents the real-valued input, $S$ the quantization scale, $r_{q}$ the quantized value, and $\hat{r}$ the dequantized value. The function $Q(\cdot)$ denotes the quantization operation and $\lfloor\cdot\rceil$ denotes rounding to the nearest integer. $Q_{N}$ and $Q_{P}$ define of the minimum and maximum quantized absolute values, respectively, ensuring the clipped value remains within the designated range. The gradient of the quantization error with respect to the scale $S$ is:
@@
These two key distinctions also offer two major benefits: 1) Learnable per-channel scales enable the model to capture and address quantization errors arising from various computations during training, including neuron dynamics like the decay process. This results in improved accuracy, as demonstrated in Section VI. This end-to-end training process ensures that, once trained, the model can be deployed directly without any additional precision loss. In contrast, statistical methods require a post-training process, and since the scales are based on weight statistics, applying them directly to biases and neuronal thresholds can cause mismatches and lead to accuracy degradation. 2) By sharing the quantization scale across weights within the same channel, along with the corresponding bias and neuronal threshold, our approach ensures consistent scaling across parameters. This eliminates the need for rescaling during weight accumulation, bias addition, and threshold comparison, enabling efficient integer-only inference. Additionally, due to the unique properties of spiking neurons, there is no need to rescale per-channel activations during inference compared to ANN quantization, as the per-channel quantization scales are absorbed into the membrane thresholds. Based on these characteristics, our hardware accelerator can perform all computations using integer arithmetic, simplifying the computational logic and memory footprint.
@@ IV-A Dataflow and Parallelism Schemes
However, a key distinction arises between FireFly-S and its predecessors. While both FireFly and FireFly v2 utilize overlay architectures, relying on off-chip memory for feature map and weight storage, FireFly-S adopts a spatial architecture to enable full inter-layer pipelining. This shift necessitates a departure from FireFly v2’s dataflow strategy. In detail, specific solutions are required to address the two main challenges posed by spatial architectures: 1) the tight on-chip memory constraints, which demand efficient memory utilization, and 2) the need for a consistent dataflow pattern to ensure efficient processing across layers and avoid pipeline stalls.
Directly applying FireFly v2’s dataflow to a spatial architecture would prove inefficient. FireFly v2, designed with off-chip memory reliance, employs an output-channel outermost loop. This approach, while functional, leads to excessive on-chip memory consumption in a spatial architecture. Computing each output channel requires access to all input feature maps, necessitating storage of all layer feature maps on-chip, amounting to $L\times F_{h_{o}}\times F_{w_{o}}\times T\times C_{i}$ memory, where $L$ represents the number of layers. Such demands would exceed the limits of on-chip memory.
Besides, the output channel in this layer is the input channel in the next layer, so it is challenging to design a consistent dataflow pattern. But we can minimize the storage and data movement. Employing an output-channel outermost loop and an input-channel innermost loop, as in FireFly v2, would lead to significant pipeline stalls in a spatial architecture. Each layer would need to wait for all output channel computations of the previous layer to complete, consequently, it could not match the need of non-stall pipeline for spatial architecture.
@@ IV-B Sparse Representation Analysis
This analysis, conducted within a specific experimental environment, simulating a convolutional operation with parameters $(T,C_{o},C_{i},F_{h_{o}},F_{w_{o}},K_{h},K_{w})=(4,64,32,16,16,3,3)$. The setup represents a single layer within the 9-layer network for CIFAR-10 recognition discussed in Section VI, and the selection does not affect the overall observed trend. Fig. 4 illustrates the throughput-to-storage ratio across various sparsity levels, comparing the performance of COO, CSR, and Bitmap sparse representations against a dense representation. While COO and CSR outperform Bitmap at extremely high sparsity levels (above 95%) due to Bitmap’s larger memory footprint, Bitmap consistently demonstrates superior performance at other sparsity levels. This advantage is further amplified with larger vector pair lengths, extending its applicability to a wider range of sparsity. Notably, in Figure 4, the sparsity level is shared by both synaptic weights and activation spikes. Experiments indicate that the sparsity of activation spikes ranges from 70% to 90%. Therefore, the sparsity of the weights must be significantly higher to reach the threshold where Bitmap becomes less efficient than COO or CSR. Given these observations, we adopt the Bitmap representation in FireFly-S, prioritizing its greater generality and higher throughput-to-storage ratio across a wide range of sparsity levels.
@@ V Hardware Micro Architecture
The hardware microarchitecture of FireFly-S is intricately designed to harness the dual-side sparsity characteristic of SNNs. It features a spatial layout that incorporates inter-layer pipelining, as illustrated in Fig. 5. Each pipeline stage is specifically designed to handle a single layer of inference within the FPGA chip, equipped with on-chip parameters and specialized modules including Padding, Dataflow Orchestrator, Dual-side Sparsity Detector, and Maxpooling.
@@
The Padding module dynamically adjusts input dimensions for convolutional layers, adding padding in ”same” mode to match output and input sizes, while bypassing padding for ”valid” mode or fully connected layers. The Dataflow Orchestrator is pivotal in ensuring consistent dataflow between layers to maintain pipeline efficiency. It also efficiently manages the im2col operation for SNNs, essential for reorganizing input data into a format suitable for convolution operations, all while maintaining a membrane potential storage-free approach. The Dual-side Sparsity Detector module optimizes computations by identifying matching non-zero elements in both activation spikes and synaptic weights, enabling computational savings by bypassing unnecessary operations. Then, it generates specific read commands for relevant data (weights, biases, etc.) and efficiently updates neuron dynamics in the IF/LIF neuron model. Finally, the optional Maxpooling module provides spatial downsampling of the output, the output stream is then pipelined and fed into the next layer for further processing.